OpenStudy (jarp0120):

calculus:rectilinear motion an object is thrown vertically upward from the ground reaches a certain height after 2 sec and returns to the same height on descent, 8 sec later. find its initial velocity and the height in question.

OpenStudy (jarp0120):



OpenStudy (jarp0120):

is this right?

Miracrown (miracrown):

The upward line is 2s, and the total curve at the top is 8s - That seems right :)

Miracrown (miracrown):

We want to use the idea that acceleration due to gravity is constant to help us Do you know if your class prefers feet or meters? We can use either units, but I don't want to do it in a way you haven't seen =/

Miracrown (miracrown):

@jarp0120

OpenStudy (jarp0120):

feet

OpenStudy (jarp0120):

and we need to use derivative to find the answer

Miracrown (miracrown):

So an important general law all airborne objects follow is this: d^2 y/dt^2 = -32.2 ft/s^2

Miracrown (miracrown):

You know the second derivative of height is a constant

OpenStudy (jarp0120):

so that's that acceleration?

Miracrown (miracrown):

Lets think mostly about the 8 seconds it was highest up first. What do you think the first derivative of its height was when it was at the very top?

Miracrown (miracrown):

The first derivative is velocity. The second derivative is acceleration We know the second derivative is constant

OpenStudy (jarp0120):

oh yes okay

Miracrown (miracrown):

Can you answer my question?

Miracrown (miracrown):

If youre not sure yet, that's ok too, its just kind of a shortcut if you realize it

OpenStudy (jarp0120):

i don't know

Miracrown (miracrown):

That's ok, lets move forward with the problem a bit and readdress it later, when it comes up. So we have this constant second derivative...But that's not very descriptive. Can you use it to get an equation for the first derivative?

OpenStudy (jarp0120):

is the first derivative v=vo-32t?

Miracrown (miracrown):

Yes, that's very well put. :)

Miracrown (miracrown):

So, do you have any ideas how we might find vo? Its ok if you dont, but if you do they could be useful

OpenStudy (jarp0120):

i don't really know

Miracrown (miracrown):

Dont worry, even if you cant see a way yet, we can forge ahead blindly and well find something useful eventually :P

OpenStudy (jarp0120):

thank you

Miracrown (miracrown):

So often when unsure the best thing to do in real life is just fight forward. So lets do that. We found an equation for velocity. Can you find one for height? (y, or the 0th derivative)

OpenStudy (jarp0120):

s=vo-16t^2?

OpenStudy (jarp0120):

s= Vot-16t^2?

OpenStudy (jarp0120):

@miracrown?

OpenStudy (jarp0120):

is the length in 2 sec = to the length in 8 sec?

Miracrown (miracrown):

@jarp0120 you said before: s= Vot-16t^2? That's technically correct...did you have a d0 at first, and replace it with zero? or an s0, I guess

OpenStudy (jarp0120):